OpenStudy (annie12m):

MEDAL + FAN What is the solution to the equation? 7(5 + 9) = 7 • x + 7 • 9 A. x = 1 B. x = 5 C. x = 7 D. x = 9

OpenStudy (anonymous):

try substituting your choices to your equation

OpenStudy (annie12m):

i tried but i keep on getting -23, but thats not an optional choice :(

RhondaSommer (rhondasommer):

i just helped you :(

OpenStudy (annie12m):

i didnt get it

RhondaSommer (rhondasommer):

okay

OpenStudy (anonymous):

you have to see if they are equal

RhondaSommer (rhondasommer):

7 * x + 7 * 9 can be simplified to what

RhondaSommer (rhondasommer):

7 is on the outside of the parenthesis

OpenStudy (annie12m):

7x + 7 times 9 equals 7x + 63 equals x = 9

RhondaSommer (rhondasommer):

no; 7(x+9) is what you're looking for in this equation

RhondaSommer (rhondasommer):

your "undisributing"

OpenStudy (annie12m):

7x + 63 x = 9

RhondaSommer (rhondasommer):

no.

RhondaSommer (rhondasommer):

stop

RhondaSommer (rhondasommer):

look at what i sent you. 7(x+9)

RhondaSommer (rhondasommer):

7(x+9) = 7(5+9)

RhondaSommer (rhondasommer):

lok at it for a second and tell me the answer

RhondaSommer (rhondasommer):

look*

OpenStudy (annie12m):

yes cause you divide 7 from 7x to get 1x which x by its self equals x and then you do it to the other side which is 63 divided by 7 to get 9 which is x = 9

RhondaSommer (rhondasommer):

no.

RhondaSommer (rhondasommer):

look

RhondaSommer (rhondasommer):

even if you simplify

OpenStudy (annie12m):

oh oh wait

RhondaSommer (rhondasommer):

look hold on.

RhondaSommer (rhondasommer):

7(x+9) = 7(5+9) 7x + 63 = 35 + 63 7x = 35 (subtract 63 on both sides) x = 5 (divide 7 on both sides)

RhondaSommer (rhondasommer):

or

OpenStudy (annie12m):

sorry i forgot its like this 7(x+9) ok so you distribute like 7x + 63 and then you cant do anything else cause its on the same side

RhondaSommer (rhondasommer):

7(x+9) = 7(5+9) x+9 = 5+9 (divide 7) x = 5 (subtract 9 on both sides)

OpenStudy (annie12m):

7 (5+9) = 7 * 5 + 7 * 9 35 + 63 = 35 + 63 then they equal the same so your right, im sorry

RhondaSommer (rhondasommer):

its okay :)

RhondaSommer (rhondasommer):

you just misunderstood

OpenStudy (annie12m):

ya i know what i done wrong, thanks for the help

OpenStudy (annie12m):

can you help me with this one now?

OpenStudy (annie12m):

@RhondaSommer

OpenStudy (annie12m):

@RhondaSommer please help!

OpenStudy (annie12m):

@safyousuff97